On September 1, 2018, Speaker Company purchased equipment for $66,000, and installation costs totaled $5,000. The equipment has an estimated useful life of 8 years and an estimated salvage value of $3,000. If the company uses the straight line method, the depreciation expense for 2018 would be:

Select one:

a. $2,625

b. $8,500

c. $2,833.33

d. $4,250

Bren Company purchased a patent for $36,000. The patent is expected to have a finite life of 10 years even though its legal life is 17 years. The amortization for the first year is:

Select one:

a. None of these.

b. $36,000.

c. $1,800

d. $3,600.

e. $3,240.

f. $2,118.

A fixed asset with a cost of $15,000 and accumulated depreciation of $12,500 is sold for $1,750. What is the amount of gain or loss on disposal of the fixed asset?

Select one:

a. $750 loss

b. $2,500 loss

c. $750 gain

d. $2,500 gain

A patent was purchased for $585,000 with a legal life of 20 years. Management estimates that the patent has a 12-year economic life. The annual amortization expense will be:

Select one:

a. $48,750.

b. $0 because patents cannot be amortized

c. $29,250.

d. $585,000.
Answer #1

1)
Depreciation expense = ($66,000 + $5,000 - $3,000) / 8 = $8,500

Option b

2)

Patents should be amortised over finite life.
Amortisation = $36,000 / 10 = $3,600

Option d

3)

Book value of asset = $15,000 - $12,500 = $2,500

Asset is sold for $1,750 which is less than book value.
Loss = $2,500 - $1,750 = $750

Option a

4)

Amortisation should be made over economic life.
Amortisation = $585,000 / 12 = $48,750

Option a